OpenStudy (anonymous):

Rebecca deposited $5,338.18 into a savings account with an interest rate of 5.9% compounded twice a year. About how long will it take for the account to be worth $10,000?

OpenStudy (ghostgate):

Well I believe first of all you will have to take $5,338.18 and find the 5.9% value of that which is $5,338.18 x 0.05 = 5.9% which is $267 estimated. But if its twice a year than add that so: $5,605.089 And now find 5.9% of that which should be: Well actually before I continue is this the same 5.9% value each year or does the 5.9% value raise each time the interest is added?

OpenStudy (anonymous):

no it's the same

OpenStudy (ghostgate):

So $267 dollars(estimated) would be your constant increase.

OpenStudy (ghostgate):

Well if so than it will basically be $10,000 - $5,338.18 = $4,661.82 So $4,661.82 divided by 266.909 = 17.5 times(estimated) Therefore that divided by 2 = 8.75 Therefore it will take 8.75 years(if the interest remains the same) to reach $10,000 dollars. I just hope my math is correct. Anyways hope that helps! G'day!

OpenStudy (anonymous):

a. 15 years, 11 months b. 32 years, 5 months c. 10 years, 10 months d. 11 years, 0 months >_<

OpenStudy (ghostgate):

Hmmm sorry. Okay lets see.

OpenStudy (anonymous):

No problem! ^o^

OpenStudy (ghostgate):

Rebecca deposited $5,338.18 into a savings account with an interest rate of 5.9% compounded twice a year. About how long will it take for the account to be worth $10,000? So this is our question. $5,338.18 dollars with a 5.9% interest rate(that stays the same) and it adds on twice a year and the total is supposed to equal $10,000 dollars right?

OpenStudy (anonymous):

Yes

OpenStudy (ghostgate):

Alright this will take awhile: I'm going to round to the thousandths cause I don't like typing huge numbers. ____#1____|__5.9%__|5.9% of #1|_#1__|Sum of 5.9% and #1|__Amount of Years__| 5,338.18 x 0.059 = 314.953 + (#1) = 5,663.133 = Half a Year. 5,663.133 x 0.059 = 333.535 + (#1) = 5,986.667 = A full year. 5,986.667 x 0.059 = 353.213 + (#1) = 6,339.881 = 1.5 years. 6,339.881 x 0.059 = 374.053 + (#1) = 6,713.934 = 2 years. 6,713.934 x 0.059 = 396.122 + (#1) = 7,110.056 = 2.5 years. 7,110.056 x 0.059 = 419.493 + (#1) = 7,529.549 = 3 years. This is what I have so far.

OpenStudy (ghostgate):

I believe I'm doing it correctly, but I may come up with the wrong answer.

OpenStudy (anonymous):

I think you are doing it right now im not sure but here is the formula A=P(1+r/n)^nt

OpenStudy (ghostgate):

That would be: Amount = Percentage(1 + rate divided by number)^number x time?

OpenStudy (ghostgate):

So: The amount would be $10,000 right? The Percentage is 5.9% The Rate is twice a year so 2 The Number is 5,338.18 Number again is it above. Time = 0.5 of a year. And we plug in these numbers right?

OpenStudy (ghostgate):

Okay I'm sorry I don't know where my brain is today. I just can't seem to figure it out for you sorry.

OpenStudy (ghostgate):

Wait my brain is rethinking this. A = P(1 + r/n)^nt So: 10,000 = $5,338.18(1 + 0.059 / 2)^2 x t 10,000 = $5,338.18 x (1.0295)2t So We can't multiply 1.0295 and 2t together cause there not like terms so lets multiply: 5,338.18 x 1.0295 together which equals = 5,495.65631 Therefore since the equation is now: 10,000 = 5,495.65631 + 2t we do this: 10,000 - 5,495.65631 = 5,495.65631 - (5,495.etc) + 2t This will equal: 4,504.34369 = 2t 4,504.34369 divided by 2 = t Equals: 2,252.##### Okay what the crap? I did everything correctly right? *sighs* I hope that helps you get an understanding of what to do.